OpenStudy (anonymous):

State how many imaginary and real zeros the function has. f(x) = x^3 - 20x^2 + 123x - 216

ganeshie8 (ganeshie8):

familiar wid Descartes rule of signs ?

OpenStudy (anonymous):

I am familiar with it but need to be reminded

ganeshie8 (ganeshie8):

ok take f(x) how many sign changes do u see ?

OpenStudy (anonymous):

2?

ganeshie8 (ganeshie8):

try again

ganeshie8 (ganeshie8):

u need to see how many times the sign is changing

ganeshie8 (ganeshie8):

f(x) = x^3 - 20x^2 + 123x - 216 ^ ^ ^

OpenStudy (anonymous):

Okay 3

ganeshie8 (ganeshie8):

good :), now find f(-x), and count sign changes

ganeshie8 (ganeshie8):

f(x) = x^3 - 20x^2 + 123x - 216 f(-x) = ?

OpenStudy (anonymous):

So do the inverse? I'm stupid forgive me.

ganeshie8 (ganeshie8):

hey first tim eyou doing this, i can understand...

ganeshie8 (ganeshie8):

to find f(-x), just replace all x with -x

ganeshie8 (ganeshie8):

f(x) = x^3 - 20x^2 + 123x - 216 f(-x) = (-x)^2 - 20(-x)^2 + 123(-x) - 216

OpenStudy (anonymous):

x-20x-123-216 or am I completely off

ganeshie8 (ganeshie8):

you're right ! but i was wrong

ganeshie8 (ganeshie8):

f(x) = x^3 - 20x^2 + 123x - 216 f(-x) = (-x)^3 - 20(-x)^2 + 123(-x) - 216

ganeshie8 (ganeshie8):

i did a typo, its (-x)^3, its not (-x)^2

ganeshie8 (ganeshie8):

f(x) = x^3 - 20x^2 + 123x - 216 f(-x) = (-x)^3 - 20(-x)^2 + 123(-x) - 216 = -x^3 - 20x^2 - 123x - 216

ganeshie8 (ganeshie8):

now tell me, how many sign changes u see in f(-x) ?

OpenStudy (anonymous):

None?

ganeshie8 (ganeshie8):

Yes, so we conclude like this :- 3 sign changes in f(x), so number of positive zeroes = 3, or 1 0 sign changes in f(x), so number of negative zeroes = 0 so total number of real zeroes = 3 or 1

ganeshie8 (ganeshie8):

that gives number of imaginary zeroes = 0 or 2

ganeshie8 (ganeshie8):

see if that makes some sense

OpenStudy (anonymous):

0 imaginary and 3 real?

ganeshie8 (ganeshie8):

0 imaginary and 3 real or 2 imaginary and 1 real

ganeshie8 (ganeshie8):

Both are true.

ganeshie8 (ganeshie8):

Both are true according to Descarted rule of signs. But when we solve it we will see that "0 imaginary and 3 real" is the correct one.

OpenStudy (anonymous):

Thank you so much, you are a champion.

ganeshie8 (ganeshie8):

omg thank you :) im not champion, you're the champion cuz you learning things very fast :)